# Feature Request Analysis and Processing: Tasks, Techniques, and Trends

### v2

@@ 8. Threats to Validity
External Validity. The scope of the included studies constrains the generalizability of our findings. This review specifically targets research involving feature requests within the context of software engineering practices. Broader studies on user feedback or those situated outside the software engineering domain were excluded. Consequently, the applicability of our results may be limited when extended to other forms of user-centered requirements engineering.
@@ 9. Conclusion
User feature requests, sourced directly from end-users, are increasingly vital for software evolution but pose unique challenges due to their volume, informality, and divergence from traditional requirements. This paper presents a comprehensive survey of 140 studies, synthesizing the current research landscape on analyzing and processing these requests throughout the software lifecycle. We examine techniques for twelve research topics on feature requests, organizing findings within a requirements engineering framework. The survey concludes by highlighting significant research gaps and future directions.
@@ Acknowledgements.
To Robert, for the bagels and explaining CMYK and color spaces.

### v3

@@ 8. Threats to Validity
External Validity. The scope of the included studies constrains the generalizability of our findings. This review specifically targets research involving feature requests within the context of software engineering practices. Broader studies on user feedback or those situated outside the software engineering domain were excluded. Consequently, the applicability of our results may be limited when extended to other forms of user-centered RE.
@@ 9. Conclusion
User feature requests, sourced directly from end-users, are increasingly vital for software evolution but pose unique challenges due to their volume, informality, and divergence from traditional requirements. This paper presents a comprehensive survey of 140 studies, synthesizing the current research landscape on analyzing and processing these requests throughout the software lifecycle. We examine techniques for twelve research topics on feature requests, organizing findings within a RE framework. The survey concludes by highlighting significant research gaps and future directions.
@@ Acknowledgements.
This work was supported by the National Natural Science Foundation of China (61802167).